PLTL Worksheet Spring 2018 Chapter 3 & 11 Nameah laicr 1. Balance the following equations: 4 H20 + 2. Ammonia, NHs can be synthesized by the following reaction: Starting with 86.3 g NO and 25.6 g H2, find the theoretical yield of ammonia in grams. 3. Titanium metal can be obtained from its oxide according to the following balanced equation: TIO2ts) +2Gy→Ties) + 2 coug) When 28.6 kg of C is allowed to react with 88.2 kg of TiO2, 42.8 kg of Ti is produced. Find the limiting reactant, theoretical yield (in kg), and percent yield.
Answer #1

Balanced equation:
2 C7H14 + 21 O2 ===> 14 CO2 + 14 H2O

Reaction type: combustion

3 Ca(OH)2 + 2 H3PO4 = Ca3(PO4)2 + 6 H2O
Reaction type: double replacement

5H2O2 + 2KMnO4 + 3H2SO4 ===> 5O2 + 2MnSO4 + K2SO4 + 8H2O

Redoc reaction

Quesion 2

Balanced equation:
2 NO + 5 H2 ===> 2 NH3 + 2 H2O

Reaction type: double replacement

86.3 gm of NO = 86.3 /30 = 2.876 Moles

25.6 gm of H2 = 25.6 / 2.016 = 12.69 Moles

Limiting regent is NO

Molesk of NH3 to be produced =  2.876 Moles

Mass of NH3 produced (Thoretical yield) = 2.876 x 17.03 = 48.98 gm

Question 3

Balanced equation:
TiO2 + 2 C ===> 2 CO + Ti

Reaction type: single replacement

28.6 Kg of C = 28.6 x 103 / 12.01 =  2381.2 Moles

88.2 Kg of TiO2 = 88.2 / 79.86 = 1104.3 Moles  

Limiting reagent is TiO2

Moles of Ti Produced =  1104.35 Moles

Mass of TiO2 produced (Thoretical Yeld) = 1104.35 x 47.86 / 1000 =  52.862 Kg

Percentage yield = 42.8 x 100 / 52.862 = 80.965 %